OpenStudy (anonymous):

Solve the following system of linear equations algebraically for all values of x and y: -14=-20y-7x 10y+4=2x

OpenStudy (neer2890):

-7x-20y=-14 2x-10y=4 multiply 2nd equation by 2 and subtract from 1st. solve for x put the value of x in any equation and find y.

OpenStudy (anonymous):

7x+20y=14 4x-20y=8 11x=22 x=2 &y=(2*2-4)/10=0 so x=2 &y=0